OpenStudy (anonymous):

how would you do 2+1/4 divided by 3/8 in order of operations

OpenStudy (radar):

Order of operations you have addition and division which do you think should be done first?

OpenStudy (anonymous):

I Know its division first and I also flipped the second # but i got a totally different answer then my book

OpenStudy (radar):

OK some clarfication on my part is the 3/8 to be divided only into the 1/4 or is the problem (2+1/4)/3/8?

OpenStudy (anonymous):

There is no parenthesis it actually is 2+1/4 over 3/8

OpenStudy (radar):

division with fraction is as you said flip and multiply. 1/4 *8/3=8/12=2/3 then add the 2 should give you 2 2/3

OpenStudy (anonymous):

thats exactly what i got but the answer in the back of the book is 6

OpenStudy (radar):

Ok it is divided in the total thin (2+1/4)/3/8 It was stated in a very unclear manner. The 3/8 is to be divided into the quantity 2 + 1/4. Why they did not use parens puzzles me, It seems like they are tryuing to trick you.

OpenStudy (anonymous):

ok thanks Radar

OpenStudy (radar):

First add the 2 + 1/4 getting 9/4 then 9/4 / 3/8= 9/4*8/3=6

OpenStudy (radar):

Good luck. I am gonna take a break.